Find the critical point of the function. Then use the second derivative test to classify the nature of this point, if possible. (If an answer does not exist, enter DNE.)

f(x, y) = x2 − 4xy + 2y2 + 4x + 8y + 8

critical point

(x, y)=

classification ---Select--- :relative maximum, relative minimum ,saddle point, inconclusive ,no critical points

Finally, determine the relative extrema of the function. (If an answer does not exist, enter DNE.)

relative minimum value=

relative maximum value=
Answer #1

Given function is f(x,y) = x² - 4 xy + 2y² + 4x + 8y +8 Now, for (x,y) = 2x - 4y + 4 6 fg (2.7) = -42+44 + 8 __ fagy = 4 we k